OpenStudy (anonymous):

if f(x)=the cubed root of (x-1), g(x)=x^2+1 what is g(f(x))?

OpenStudy (anonymous):

f(x)=\[\sqrt[3]{x-1}\] g(x)= \[x ^{2}+1\] g(f(x)))=\[((\sqrt[3]{x-1})^2+1)\]

OpenStudy (anonymous):

It is possible to simplify further